# Newton polytopes in cluster algebras and $\tau$-tilting theory

### v1

@@ 3.3. Bongartz completion in $\tau$-tilting theory
A module $M$ in a subcategory $\mathcal{C}$ of $\operatorname{\mathsf{mod}}\nolimits A$ is said to be Ext-projective in $\mathcal{C}$, if $\operatorname{\mathsf{Ext}}_{A}^{1}(M,\mathcal{C})=0$. We denote by $\mathcal{P}(\mathcal{C})$ the direct sum of one copy of each of the indecomposable Ext-projective objects in $\mathcal{C}$ up to isomorphism.
@@
It is each to see that the left Bongartz completion $(M^{-},P^{-})$ of $(U,Q)$ can be characterized by the following two conditions:
@@
4. $F$-invariant and Newton polytopes of $F$-polynomials
@@ 4.1. $F$-invariant of cluster monomials
be the canonical expressions of $u$ and $v$ with repect to vertex $t\in\mathbb{T}_{n}$.
@@
In particular, the $F$-invariant $(u\mid\mid v)_{F}=F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]$ only depends on $u$ and $u^{\prime}$, not on the choice of vertex $t\in\mathbb{T}_{n}$.
@@
Let $u={\bf x}_{t}^{\bf h}=\prod_{i=1}^{n}x_{i;t}^{h_{i}}$ and $v$ be two cluster monomials of $\mathcal{A}$. Then we have $(u\mid\mid v)_{F}=\sum_{i=1}^{n}h_{i}(x_{i;t}\mid\mid v)_{F}$. Moreover, we have
@@ 4.2. Newton polytopes of $F$-polynomials
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ and $D=diag(d_{1},\ldots,d_{n})$ a fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. Let $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ be a mutation in $\mathcal{A}$. Then $(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=d_{k}$. If moreover, the $k$-th column of $C_{t}^{t_{0}}$ is non-positive, then we have
$F_{x_{k;t^{\prime}}}^{t_{0}}[{\bf g}_{x_{k;t}}]=0\quad\text{and}\quad F_{x_{k;t}}^{t_{0}}[{\bf g}_{x_{k;t^{\prime}}}]=d_{k}.$
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t}[D{\bf g}_{k;t^{\prime}}^{t}]+F_{x_{k;t^{\prime}}}^{t}[D{\bf g}_{x_{k;t}}^{t}]=0+(1+y_{k})[D{\bf e}_{k}]=d_{k}.$
@@
$d_{k}=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[D{\bf g}_{k;t^{\prime}}^{t}]+F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t}].$
Now let us prove that $F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{k;t}^{t}]=0$ under the assumption that the $k$-th column ${\bf c}_{k;t}$ of $C_{t}^{t_{0}}$ is non-positive. In this case, we have ${\bf c}_{k;t}\in\mathbb{Z}_{\geq 0}^{n}$. By the mutation rule [fomin_zelevinsky_2007, Proposition 5.1] of $F$-polynomials, we have
@@
for any ${\bf r}\in\mathbb{R}^{n}$. By Lemma 4.6, we know that $F_{x_{i;t}}^{t_{0}}[D{\bf g}_{k;t}^{t_{0}}]=0$ for any $i\in[1,n]$. Then by taking ${\bf r}=D{\bf g}_{x_{k;t}}^{t_{0}}$ in (4.1), we get
@@
Thus $F_{x_{k;t}}^{t_{0}}[{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}-F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=d_{k}-0=d_{k}$. ∎
@@
We know that $u=u^{\prime}\cdot x_{k;t}^{a_{k}}$ and $v=v^{\prime}\cdot x_{k;t}^{b_{k}}=v^{\prime}\cdot x_{k;t}^{a_{k}}$. Denote by $H=(F_{x_{k;t}}^{t_{0}})^{a_{k}}$. Then we have $F_{u}^{t_{0}}=F_{u^{\prime}}^{t_{0}}\cdot H$ and $F_{v}^{t_{0}}=F_{v^{\prime}}^{t_{0}}\cdot H$. By Proposition 2.4, we have $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{u^{\prime}}^{t_{0}})+\mathsf{P}(H)$ and $\mathsf{P}(F_{v}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})+\mathsf{P}(H)$. Since $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and by Corollary 2.2, we get $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$. ∎
@@
Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$. If $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$, then the product $uv$ remains a cluster monomials of $\mathcal{A}$.
@@
Thus to show $u=v$, it suffices to show $u^{\prime}=v^{\prime}$. Now $u^{\prime}$ and $v^{\prime}$ are two cluster monomials in non-initial cluster variables satisfying $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$ and $|\operatorname{\mathsf{supp}}(u^{\prime}v^{\prime})|<|\operatorname{\mathsf{supp}}(uv)|$. We can therefore apply the reduction and complete the proof. ∎
5. $F$-invariant and Newton polytopes of $\tau$-rigid modules
@@ 5.2. Newton polytopes of $\tau$-rigid modules
By Lemma 5.12, we know that the direct sum $U\oplus V$ is $\tau$-rigid. We will prove that $U\cong V$ by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. We can assume $|U\oplus V|>0$. Otherwise, we have $U=V=0$ and the desired result holds trivially.
@@
By Corollary 3.10, we know that $\mathsf{P}(U)=\mathsf{P}(U^{\prime})+\mathsf{P}(M_{k}^{a_{k}})$ and $\mathsf{P}(V)=\mathsf{P}(V^{\prime})+\mathsf{P}(M_{k}^{b_{k}})$. Since $\mathsf{P}(U)=\mathsf{P}(V)$ and by Corollary 2.2, we get $\mathsf{P}(U^{\prime})=\mathsf{P}(V^{\prime})$.
@@ 5.3. Newton polytopes of left finite multi-semibricks
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is left finite, if the smallest torsion class $\langle M\rangle_{\rm tors}$ containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$. A module $C\in\operatorname{\mathsf{mod}}\nolimits A$ is called a brick, if $\operatorname{\mathsf{End}}_{A}(C)\cong\mathsf{k}$. A set $\Omega$ of iso-classes of bricks is called a semibrick, if $\operatorname{\mathsf{Hom}}_{A}(C,C^{\prime})=0$ for any $C,C^{\prime}\in\Omega$ with $C\neq C^{\prime}$. A semibrick $\Omega$ is left finite, if the smallest torsion class $\langle\Omega\rangle_{\rm tors}$ containing the bricks in $\Omega$ is a functorially finite torsion class.
@@
    Denote by $\Omega:=\{C_{i}\}_{i\in I}$ the set of labeling bricks for the left mutations of $\mathcal{M}$. Then $\Omega$ is a left finite semibrick and we have $\langle\Omega\rangle_{\rm tors}=\operatorname{\mathsf{Fac}}M$.

### v2

@@ 3.3. Bongartz completion in $\tau$-tilting theory
A module $M$ in a subcategory $\mathcal{C}$ of $\operatorname{\mathsf{mod}}\nolimits A$ is said to be Ext-projective in $\mathcal{C}$, if $\operatorname{\mathsf{Ext}}_{A}^{1}(M,X)=0$ for any $X\in\mathcal{C}$. We denote by $\mathcal{P}(\mathcal{C})$ the direct sum of one copy of each of the indecomposable Ext-projective objects in $\mathcal{C}$ up to isomorphism.
@@
It is easy to see that the left Bongartz completion $(M^{-},P^{-})$ of $(U,Q)$ can be characterized by the following two conditions:
@@
4. $F$-invariant and Newton polytopes in cluster algebras
@@ 4.1. $F$-invariant of cluster monomials
be the canonical expressions of $u$ and $v$ with repect to a vertex $t\in\mathbb{T}_{n}$.
@@
In particular, the $F$-invariant $(u\mid\mid v)_{F}=F_{u}^{t}[D{\bf g}_{v}^{t}]+F_{v}^{t}[D{\bf g}_{u}^{t}]$ only depends on $u$ and $v$, not on the choice of vertex $t\in\mathbb{T}_{n}$.
@@
Let $u=\prod_{i=1}^{n}x_{i;t}^{h_{i}}$ and $v$ be two cluster monomials of $\mathcal{A}$. Then we have $(u\mid\mid v)_{F}=\sum_{i=1}^{n}h_{i}(x_{i;t}\mid\mid v)_{F}$. Moreover, we have
@@ 4.2. Newton polytopes of $F$-polynomials
Let $\mathcal{A}$ be a cluster algebra with initial seed $({\bf x}_{t_{0}},B_{t_{0}})$ and $D=diag(d_{1},\ldots,d_{n})$ a fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}$. Let $({\bf x}_{t^{\prime}},B_{t^{\prime}})=\mu_{k}({\bf x}_{t},B_{t})$ be a mutation in $\mathcal{A}$. Then
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=d_{k}.$
If moreover, the $k$-th column of $C_{t}^{t_{0}}$ is non-positive, then we have
$F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=0\quad\text{and}\quad F_{x_{k;t}}^{t_{0}}[D{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=d_{k}.$
@@
$(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t}[D{\bf g}_{x_{k;t^{\prime}}}^{t}]+F_{x_{k;t^{\prime}}}^{t}[D{\bf g}_{x_{k;t}}^{t}]=0+(1+y_{k})[D{\bf e}_{k}]=d_{k}.$
@@
$d_{k}=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}=F_{x_{k;t}}^{t_{0}}[D{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]+F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}].$
Now let us prove that $F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=0$ under the assumption that the $k$-th column ${\bf c}_{k;t}$ of $C_{t}^{t_{0}}$ is non-positive. In this case, we have ${\bf c}_{k;t}\in\mathbb{Z}_{\geq 0}^{n}$. By the mutation rule [fomin_zelevinsky_2007, Proposition 5.1] of $F$-polynomials, we have
@@
for any ${\bf r}\in\mathbb{R}^{n}$. By Lemma 4.6, we know that $F_{x_{i;t}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=0$ for any $i\in[1,n]$. Then by taking ${\bf r}=D{\bf g}_{x_{k;t}}^{t_{0}}$ in (4.1), we get
@@
Thus $F_{x_{k;t}}^{t_{0}}[D{\bf g}_{x_{k;t^{\prime}}}^{t_{0}}]=(x_{k;t}\mid\mid x_{k;t^{\prime}})_{F}-F_{x_{k;t^{\prime}}}^{t_{0}}[D{\bf g}_{x_{k;t}}^{t_{0}}]=d_{k}-0=d_{k}$. ∎
@@
We know that $u=u^{\prime}\cdot x_{k;t}^{a_{k}}$ and $v=v^{\prime}\cdot x_{k;t}^{b_{k}}=v^{\prime}\cdot x_{k;t}^{a_{k}}$. Let $H:=(F_{x_{k;t}}^{t_{0}})^{a_{k}}$. Then we know $F_{u}^{t_{0}}=F_{u^{\prime}}^{t_{0}}\cdot H$ and $F_{v}^{t_{0}}=F_{v^{\prime}}^{t_{0}}\cdot H$. By Proposition 2.4, we have $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{u^{\prime}}^{t_{0}})+\mathsf{P}(H)$ and $\mathsf{P}(F_{v}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})+\mathsf{P}(H)$. Since $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$ and by Corollary 2.2, we get $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$. ∎
@@
Let $u$ and $v$ be two cluster monomials of $\mathcal{A}$. If $\mathsf{P}(F_{u}^{t_{0}})=\mathsf{P}(F_{v}^{t_{0}})$, then the product $uv$ remains a cluster monomial of $\mathcal{A}$.
@@
Thus to show $u=v$, it suffices to show $u^{\prime}=v^{\prime}$. Now $u^{\prime}$ and $v^{\prime}$ are two cluster monomials in non-initial cluster variables satisfying $\mathsf{P}(F_{u^{\prime}}^{t_{0}})=\mathsf{P}(F_{v^{\prime}}^{t_{0}})$ and $|\operatorname{\mathsf{supp}}(u^{\prime}v^{\prime})|<|\operatorname{\mathsf{supp}}(uv)|$. Thus, by repeatedly applying the reduction argument, we complete the proof. ∎
5. $F$-invariant and Newton polytopes in $\tau$-tilting theory
@@ 5.1. $F$-invariant of decorated modules
For a module $X\in\operatorname{\mathsf{mod}}\nolimits A$, we denote by $\underline{\dim}X\in\mathbb{N}^{n}$ the dimension vector of $X$.
@@ 5.2. Newton polytopes of $\tau$-rigid modules
By Lemma 5.12, we know that the direct sum $U\oplus V$ is $\tau$-rigid. We will prove that $U\cong V$ by reducing the number $|U\oplus V|$ of iso-classes of indecomposable direct summands of $U\oplus V$. We may assume $|U\oplus V|>0$; otherwise, we have $U=V=0$ and the desired result holds trivially.
@@
By Corollary 3.11, we know that $\mathsf{P}(U)=\mathsf{P}(U^{\prime})+\mathsf{P}(M_{k}^{a_{k}})$ and $\mathsf{P}(V)=\mathsf{P}(V^{\prime})+\mathsf{P}(M_{k}^{a_{k}})$. Since $\mathsf{P}(U)=\mathsf{P}(V)$ and by Corollary 2.2, we get $\mathsf{P}(U^{\prime})=\mathsf{P}(V^{\prime})$.
@@ 5.3. Newton polytopes of left finite multi-semibricks
A module $M\in\operatorname{\mathsf{mod}}\nolimits A$ is called left finite, if the smallest torsion class $\langle M\rangle_{\rm tors}$ containing $M$ is a functorially finite torsion class in $\operatorname{\mathsf{mod}}\nolimits A$. A module $C\in\operatorname{\mathsf{mod}}\nolimits A$ is called a brick, if $\operatorname{\mathsf{End}}_{A}(C)\cong\mathsf{k}$. A set $\Omega$ of iso-classes of bricks is called a semibrick, if $\operatorname{\mathsf{Hom}}_{A}(C,C^{\prime})=0$ for any $C,C^{\prime}\in\Omega$ with $C\neq C^{\prime}$. A semibrick $\Omega$ is left finite, if the smallest torsion class $\langle\Omega\rangle_{\rm tors}$ containing the bricks in $\Omega$ is a functorially finite torsion class.
@@
    Let $\Omega:=\{C_{i}\}_{i\in I}$ be the set of labeling bricks for the left mutations of $\mathcal{M}$. Then $\Omega$ is a left finite semibrick and we have $\langle\Omega\rangle_{\rm tors}=\operatorname{\mathsf{Fac}}M$.